# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ 2 Experimental Methods
We report on measurements performed during three experimental campaigns in the water channel facility at NTNU Trondheim, shown in figure 1a. A pump system circulates water through the test section of dimensions $11.2$ m $\times$ $1.8$ m $\times$ $1.0$ m (length $\times$ width $\times$ height). An active grid at the inlet of the test section allowed the turbulence to be generated and varied. The grid consists of square wings measuring $M=10$ cm across the diagonal, attached to $18$ vertically and $10$ horizontally oriented bars, each controlled by a stepper motor. Several different active-grid actuation cases were investigated, listed in Appendix A. The grid wings were rotated with random rotational velocity, acceleration, and period within set limits (Smeltzer et al., 2023; Hearst & Lavoie, 2015), or in one case flapped back and forth between two positions at irregular time intervals. The instantaneous rotation frequency of the grid wings varied about a mean active-grid frequency $\overline{f_{G}}$ by ±$\pm 0.5\overline{f_{G}}$ with a top-hat distribution. In experiments 1 and 2 (see sections 2.1.1 and 2.1.2) a surface plate was mounted from the grid that extends downstream approximately $1$ m to dampen surface disturbances produced by the grid. A diagram of the setup is shown in figure 1, and further details are given by Jooss et al. (2021).
Figure 1: Experimental setup: (a) side view of water channel with flow from left to right, (b) top view of measurement region for the stereo PIV setup in experiment 1, (c) longitudinal view of the planar PIV setup in experiments 2 and 3. In experiment 3, a single PIV camera was used.
A plunger wavemaker at the downstream end of the test section ($10.2$ m from the active grid) was used to generate waves propagating upstream on the current. Waves with a group velocity lower than the mean flow were unable to propagate upstream, thus preventing high-frequency wave noise and unwanted free harmonics (parasitic waves) from the wavemaker from entering the test section.
@@ 2.1 Experimental campaigns
Three separate experimental campaigns were conducted between 2020 and 2023. Experiment 1 was also reported in Smeltzer et al. (2023), where the focus was on the change in turbulent enstrophy and wave scattering. We define our lab-system coordinates so that the waves propagate in the positive $x$-direction while the direction of the mean flow (the streamwise direction) is towards negative $x$.
Experiment Cases Description
1 1.A1.D Repeated wavegroups, measurements of flow before and after.
2 2.A, 2.B Regular waves measured at low frequency for a long time.
3 3.A, 3.B Regular waves measured at high frequency in repeated bursts.
Table 1: Comparison of experimental designs. See Table 2 for experimental parameters.
@@ 2.1.1 Experiment 1: Wave groups
Cases 1.A-1.D in tables 2 and 3 are from Experiment 1. Wave groups were generated that propagated upstream atop the turbulent flows. The water depth $h$ was $0.4$ m. The velocity field was measured using stereoscopic particle image velocimetry (SPIV), measuring all three velocity components in a plane perpendicular to the mean flow located a distance $8.3883.8M=$ 8.38 m downstream from the active grid. Two 25-megapixel cameras were mounted on either side of the test section, viewing the field of view at ±$\pm 45^{\circ}$ to the $x$-axis as shown in figure 1b). The field of view was $0.12\times 0.14$ m. Particle images recorded by the two cameras were processed using a final pass of $48\times 48$ pixel interrogation window and a 50% overlap, resulting in a velocity vector spacing of about $0.8$ mm. The free surface intersection with the SPIV plane was detected from laser-induced fluorescence (LIF) images recorded by a camera viewing the plane at an oblique angle from the air side. A small amount of rhodamine-6G was added to the water generating image contrast between the air and water regions, and the free surface was detected from the image intensity gradient. Further details can be found in Smeltzer et al. (2023).
For each wave group, SPIV/LIF measurements taken during three time intervals were used: well before the group arrived, and at the leading and trailing edges of the group envelope, referred to as intervals 1, 2, and 3, respectively, as shown in figure 2b). We consider the difference in mean velocity between intervals 1 and 3 here, with interval 2 as a check to verify that the change is indeed due to waves. Values for all three intervals in all cases can be found in Supplementary Materials. The duration of the measurements for each interval, $T_{\mathrm{PIV}}$, was $10$ s, sampled at $f_{\mathrm{ac}}=8$ Hz. After each group, residual waves from reflections were allowed to dissipate for approximately five minutes before the next wave group was generated. The above procedure was performed a total of $N_{\mathrm{ens}}=60$ times to produce ensemble statistics, except for the case 1.C.2 which was performed $20$ times. In case 1.B, only vertically orientated grid bars were actuated. For case 1.A, the grid was stationary with the wings aligned with the flow in the position of least blockage (see also Appendix A). The experimental conditions are listed in table 2.
Case $h$ PIV PIV $\tau_{\mathrm{wm}}$ $T_{\mathrm{wm}}$ $T_{\mathrm{PIV}}$ $N_{\mathrm{ens}}$ $f_{\mathrm{ac}}$
$(\mathrm{m})$ plane type $(\mathrm{s})$ $(\mathrm{s})$ $(\mathrm{s})$ $(\mathrm{Hz})$
1.A 0.40 $yz$ Stereo 6 24 10 60 8
1.B 0.40 $yz$ Stereo 6 24 10 60 8
1.C.1 0.40 $yz$ Stereo 6 24 10 60 8
1.C.2 0.40 $yz$ Stereo 6 24 10 20 8
1.D 0.40 $yz$ Stereo 6 24 10 60 8
2.A 0.80 $zx$ Planar $\infty$ - 2324 1 0.86
2.B 0.80 $zx$ Planar $\infty$ - 2324 1 0.86
3.A 0.50 $zx$ Planar $\infty$ - 50 32 15
3.B 0.50 $zx$ Planar $\infty$ - 50 32 15
Table 2: Test case design parameters. $\tau_{\mathrm{wm}}$: temporal group length; $T_{\mathrm{wm}}$: duration of wavemaker actuation; $T_{\mathrm{PIV}}$: length of each acquisition interval; $N_{\mathrm{ens}}$: number of acquisition intervals; $f_{\mathrm{ac}}$: frequency of PIV acquisition. For cases 2.A3.B, $L_{\mathrm{FOV}}=8.5$ m as defined in text. Active-grid settings for each case are found in Appendix A.
@@
for $0\leq t\leq T_{\mathrm{wm}}$, where $S_{0}$ is the peak stroke, $\tau_{\mathrm{wm}}$ characterises the group width in time and $T_{\mathrm{wm}}$ was the duration over which the wavemaker plungers were actuated. The surface elevation for one wave group measured at the SPIV measurement location is shown in figure 2(a). LIF and SPIV images were acquired at a frequency $f_{\mathrm{ac}}=8$ Hz.
Figure 2: (a) Example surface elevation of a single wave group, measured by a wave probe at the measurement location. (b) Example of an ensemble-average group surface elevation amplitude envelope as a function of time normalised by the group length $\tau$ for case 1.D(see (4)). The time intervals for SPIV measurement (1-3) are shown as shaded. (c) Surface elevation measurements of one ensemble from experiment 3 (cases 3.A and 3.B), which shows the onset of a regular wave train. The red box indicates the interval used for analysis.
@@ 2.1.2 Experiment 2: Regular waves
Cases 2.A and 2.B in tables 2 and 3 are from Experiment 2. Regular waves were generated propagating upstream atop two different flows with comparable mean velocity but different levels of turbulence as controlled by the active grid. A planar PIV setup with a light sheet, orientated in the streamwise-vertical ($xz$) plane, measured the in-plane streamwise and vertical velocity components. The field of view was centred $L_{\mathrm{FOV}}=8.5$ m downstream of the surface plate’s trailing edge (see figure 1). Three cameras stacked vertically (from the top to bottom: two 16 megapixel cameras and one 5.5 megapixel camera) covered a field of view extending over the entire water depth ($h=0.80$ m) of the channel, roughly $21.7$ cm wide. The parameters for the test cases 2.A and 2.B are shown in table 2. For both flow cases, waves of two frequencies $0.94$ and $1.16$ Hz each with three different steepness values were generated. For each case, $2000$ PIV images were acquired at a sampling frequency of $0.86$ Hz; thus the total measurement time $T_{\mathrm{PIV}}$ was approximately $39$ min. Using a final pass with $48\times 48$ pixel interrogation window and a 50% overlap resulted in a velocity vector spacing of about $1.6$ mm. For both cases 2.A and 2.B, PIV measurements were also performed without wave generation to characterise ambient flow conditions, reported in Table 3.
@@ 2.1.3 Experiment 3: Onset of regular waves
Cases 3.A and 3.B in tables 2 and 3 are from Experiment 3. This set of measurements was taken at the highest acquisition frequency in our study ($15$ Hz) during a $25$ s interval just after the first arrival of regular waves for a total of $32$ times per case to obtain statistics. The water depth was $0.50$ m, and a slower mean flow of $U_{0}=0.19$ m/s was used compared to the other cases (see table 3). The two cases are separated by the different sequences for the active grid. Case 3.A is similar to case 1.B as only vertically orientated grid bars were rotated at a random, top-hat-distributed rotation speed. Similarly for case 3.B, only the vertical grid bars were actuated, however the grid bars were not rotating, but flapping over an angle of ±$\pm 60^{\circ}$ about the fully open position during random time intervals of a top-hat-distributed duration between $0.5$ and $1.0$ seconds. (A full list of experimental conditions is found in appendix A.) No surface plate downstream of the active grid was used in this experiment, and the field of view in the streamwise-vertical plane (see figure 1c) was centred a distance $L_{\mathrm{FOV}}=8.5$ m downstream of the grid.
The wavemaker was actuated at $f_{0}=1.4$ Hz. Here, two steepnesses for each flow case are investigated, yielding four different steepness-turbulence combinations. For each of the four combinations, $32$ ensembles are measured, where the wave probe measurements of one ensemble are shown in figure 2(c). In order to avoid wave breaking at the leading edge of the wave train, the wavemaker stroke is linearly ramped up to the set amplitude over a period of $6$ seconds.
For the PIV measurements, a single 25-megapixel camera was used to cover the entire water column. Using a $64\times 64$ pixel interrogation window and a 50% overlap resulted in a final velocity vector spacing of about $3$ mm. The field of view is $0.45\,\mathrm{m}\times 0.5$ m. PIV images were acquired at $f_{\mathrm{ac}}=15$ Hz. Also here, PIV measurements were performed without wave generation to characterise the ambient flow conditions reported in table 3. For these background measurements $2000$ snapshots were captured at $1.0$ Hz.
@@ 3.1 Flow and wave characteristics
The measured physical characteristics of mean flow, turbulence and waves are listed in table 3 together with some derived quantities we will make use of.
Case $U_{0}$ $k_{0}$ $[u_{\infty},v_{\infty},w_{\infty}]$ $e$ $L_{x}^{x}$ $\epsilon_{0}$ $\tau_{0}$
$(\mathrm{m/s})$ $(\mathrm{rad/m})$ $(\mathrm{cm/s})$ $(\mathrm{cm^{2}/s^{2}})$ $(\mathrm{cm})$ $(\mathrm{s})$
1.A 0.34 9.5 $[0.71,0.68,0.58]$ 0.65 $5.1$ $0.20$ 2.4
1.B 0.33 9.2 $[1.1,1.0,0.73]$ 1.3 $26$ $0.20$ 2.6
1.C 0.33 8.9,9.0 $[1.6,1.2,0.92]$ 2.4 $32$ $0.15,0.22$ 2.9
1.D 0.34 9.3 $[1.7,1.7,1.3]$ 3.7 $20$ $0.22$ 2.8
2.A.1 0.30 6.1 $[2.5,-,1.8]$ 6.4 - $0.09,0.14,0.18$
2.A.2 0.30 12.1 $[2.5,-,1.8]$ 6.4 - $0.15,0.19,0.21$
2.B.1 0.30 6.1 $[1.6,-,1.2]$ 2.7 - $0.07,0.12,0.17$
2.B.2 0.30 12.0 $[1.6,-,1.2]$ 2.7 - $0.14,0.17,0.21$
3.A 0.19 12.9 $[1.2,-,0.84]$ 1.4 5.4 $0.11,0.18$
3.B 0.19 13.0 $[0.87,-,0.70]$ 0.87 5.1 $0.11,0.18$
Case $\lambda_{0}$ $c(k_{0})$ $u_{s}(0)$ $\varpi_{0}$ $T_{\mathrm{int}}$ $T_{\mathrm{int}}/\varpi_{0}$ $|u_{\mathrm{rf}}|$ $\beta_{\mathrm{f}}(0)$
@@
1.A $0.68$ $1.0$ $4.1$ $0.65$ $4.3$ $6.6$ $3.3$
1.B $0.70$ $1.0$ $4.2$ $0.66$ $4.6$ $7.0$ $3.5$
1.C $0.71,0.70$ $1.0$ $2.3,5.1$ $0.67$ $5.1$ $7.6$ $2.2,4.7$
1.D $0.68$ $1.0$ $5.0$ $0.66$ $5.0$ $7.6$ $4.6$
2.A.1 $1.03$ $1.3$ $1.0,2.5,4.1$ $0.81$ $29.2$ $35.7$ $1.1,2.5,4.2$ $3.6,8.6,14.2$
2.A.2 $0.52$ $0.90$ $2.0,3.3,4.0$ $0.58$ $29.2$ $50.6$ $1.0,1.7,2.1$ $13.9,22.3,27.2$
2.B.1 $1.03$ $1.3$ $0.62,1.8,3.7$ $0.81$ $29.2$ $35.9$ $0.6,1.9,3.8$ $2.1,6.3,12.7$
2.B.2 $0.52$ $0.90$ $1.8,2.6,4.0$ $0.58$ $29.2$ $50.4$ $0.9,1.4,2.1$ $12.1,17.8,27.1$
3.A $0.49$ $0.87$ 1.1, 2.8 0.56 $46.1$ $82.5$ $0.8,2.2$ $12.2,32.6$
3.B $0.48$ $0.87$ 1.1, 2.8 0.56 $46.1$ $82.8$ $0.8,2.2$ $12.2,32.7$
Table 3: Measured current, turbulence and wave parameters. For cases 1.A1.D the peak steepness $\epsilon_{0\mathrm{p}}$ is reported as the value for $\epsilon_{0}$. Where several values of $\epsilon_{0}$ are listed (cases 1.C and 2.A3.B), these are referred to elsewhere as 1.C.1, 1.C.2, 2.A1.1, 2.A1.2, etc. For cases 1.A1.D, $T_{\mathrm{int}}=\sqrt{\pi}\tau_{0}$ is used, while for cases 2.A3.B, $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$; $u_{\mathrm{rf}}$ is estimated using (8).
An assumption of deep water is well satisfied since in all our cases since $k_{0}h\gtrsim 3.6$. Because deep-water waves could be assumed, $k_{0}$ and the wavemaker carrier frequency $f_{0}$ are related approximately by $2\pi f_{0}=\sqrt{gk_{0}}-U_{0}k_{0}$. Here and henceforth, $U_{0}$ is the mean absolute surface velocity (i.e., speed) in the absence of waves (i.e., the mean surface velocity is $\mathbf{U}_{0}=(-U_{0},0,0)$). The measured value of $k_{0}$ and associated wavelength $\lambda_{0}=2\pi/k_{0}$ were used in further analysis, given in table 3.
The root-mean-square (RMS) of the turbulent velocity fluctuation after subtracting the average is defined as ${\boldsymbol{u}}_{\infty}=[u_{\infty},v_{\infty},w_{\infty}]$ representing the streamwise, spanwise, and vertical components, respectively. The turbulent kinetic energy is as $e=\frac{1}{2}|{\boldsymbol{u}}_{\infty}|^{2}$. In cases 2.A3.B, only $u_{\infty}$ and $w_{\infty}$ are available, so we use instead $e=\frac{1}{2}(u_{\infty}^{2}+2w_{\infty}^{2})$ on the basis that $v_{\infty}$ is typically more similar to $w_{\infty}$ than $u_{\infty}$ when the two are different (see table 3 and Jooss et al., 2021). The mean velocity $U_{0}$ was calculated from averaging all streamwise velocities over the lower part of the field of view; averaging was performed over $-12$ cm $<z<-5$ cm for Experiment 1 and $-20$ cm $<z<-10$ cm for Experiments 2 and 3.

### v2

@@ 2 Experimental Methods
We report on measurements performed during three experimental campaigns in the water channel facility at NTNU Trondheim, shown in figure 1a. A pump system circulates water through the test section of dimensions $11.2$ m $\times$ $1.8$ m $\times$ $1.0$ m (length $\times$ width $\times$ height).
An active grid at the inlet of the test section allowed the turbulence to be generated and varied. The grid consists of square wings measuring $10$ cm across the diagonal, attached to $18$ vertically and $10$ horizontally oriented bars, each controlled independently by a stepper motor. Several different active-grid actuation cases were investigated, listed in Appendix A. The grid wings were rotated with random rotational velocity, acceleration, and period within set limits (Smeltzer et al., 2023; Hearst & Lavoie, 2015), or in one case flapped back and forth between two positions at irregular time intervals. The instantaneous rotation frequency of the grid wings varied about a mean active-grid frequency $\overline{f_{G}}$ by $\pm 0.5\overline{f_{G}}$ with a top-hat distribution. In experiments 1 and 2 (see sections 2.1.1 and 2.1.2) a surface plate was mounted from the grid that extends downstream approximately $1$ m to dampen surface disturbances produced by the grid. A diagram of the set-up is shown in figure 1, and further details are given by Jooss et al. (2021).
Figure 1: Experimental set-up: (a) side view of water channel with flow from left to right and field of view (FOV) indicated with a green rectangle, (b) top view of measurement region for the stereo particle image velocimetry (PIV) set-up in experiment 1, including positions of wave probes (WPs) and laser-induced fluoresence (LIF) camera for surface detection; (c) longitudinal view of the planar PIV set-up in experiments 2 and 3. Experiment 2 employed three stacked cameras as shown, whereas in experiment 3, a single PIV camera was used.
The bottom and sidewall boundary layers are thin enough not to reach the central part of the flow. They reach a maximum of approximately $20$ cm at the point where velocity measurements were made, and are thinner upstream where waves and turbulence interact, hence essentially all turbulence which affects our experimental results has been generated by the grid and not boundary layers. The upstream distance within which waves and turbulence interact is denoted $L_{\mathrm{FOV}}$ and is tabulated in Table 1.
A plunger wavemaker at the downstream end of the test section ($10.2$ m from the active grid) was used to generate waves propagating upstream on the current. Waves with a group velocity lower than the mean flow were unable to propagate upstream, thus preventing high-frequency wave noise and unwanted free harmonics (parasitic waves) from the wavemaker from entering the test section. Wave properties were obtained from surface elevation measured by a pair of resistive wave probes (HR Wallingford) near the measurement position.
@@ 2.1 Experimental campaigns
Three separate experimental campaigns were conducted between 2020 and 2023. Experiment 1 was also reported in Smeltzer et al. (2023), where the focus was on the change in turbulent enstrophy and wave scattering.
ID Cases Type of waves PIV PIV $h$ $f_{0}$ $f_{\mathrm{ac}}$ $N_{\text{ens}}$ $T_{\text{PIV}}$ $L_{\mathrm{FOV}}$
type plane (cm) (Hz) (Hz) (s) (m)
1 1.A1.D Repeated groups stereo yzyz 40 1.02 8 $60^{\dagger}$ 10 7.38
2 2.A, 2.B Regular waves planar zxzx 80 0.94, 1.16 0.86 1 2324 8.5
3 3.A, 3.B Regular waves planar zxzx 50 1.4 15 32 25 8.5
Table 1: Overview of the three experiments. Here, $h$ is the mean water depth, $f_{0}$ is the wave frequency in the laboratory frame (wavemaker frequency), $f_{\mathrm{ac}}$ is the PIV acquisition frequency, $N_{\text{ens}}$ is the number of repetitions per case, $T_{\mathrm{PIV}}$ is the duration of each acquisition interval and $L_{\mathrm{FOV}}$ is the distance the flow travels as a free stream upstream of the measurement field of view. $\dagger$: except case 1.C.2 where $N_{\text{ens}}=20$.
@@ 2.1.1 Experiment 1: Wave groups
Cases 1.A-1.D in Tables 13 are from Experiment 1. Wave groups were generated that propagated upstream atop the turbulent flows. The water depth $h$ was $0.4$ m. The velocity field was measured using stereoscopic particle image velocimetry (SPIV), measuring all three velocity components in a plane perpendicular to the mean flow located a distance $8.38$ m downstream from the active grid (i.e., $83.8$ grid units), and $L_{\mathrm{FOV}}=7.38$ m downstream of the trailing edge of the surface plate. Two 25-megapixel cameras were mounted on either side of the test section, viewing the field of view at $\pm 45^{\circ}$ to the $x$-axis as shown in figure 1b). The field of view was $0.12\times 0.14$ m2. Particle images recorded by the two cameras were processed using a final pass of $48\times 48$ pixel interrogation window and a 50% overlap, resulting in a velocity vector spacing of approximately $0.8$ mm. The free surface intersection with the SPIV plane was detected from laser-induced fluorescence (LIF) images recorded by a camera viewing the plane at an oblique angle from the air side. A small amount of rhodamine-6G was added to the water generating image contrast between the air and water regions, and the free surface was detected from the image intensity gradient. Further details can be found in Smeltzer et al. (2023).
For each wave group, SPIV/LIF measurements taken during three time intervals were used: well before the group arrived, and at the leading and trailing edges of the group envelope, referred to as intervals 1, 2, and 3, respectively, as shown in figure 2b). We consider the difference in mean velocity between intervals 1 and 3 here, with interval 2 as a check to verify that the change is indeed due to waves. Values for all three intervals in all cases can be found in Supplementary Materials. The duration of the measurements for each interval, $T_{\mathrm{PIV}}$, was $10$ s and PIV and LIF were sampled at $f_{\mathrm{ac}}=8$ Hz. After each group, residual waves from reflections were allowed to dissipate for approximately five minutes before the next wave group was generated. The above procedure was performed a total of $N_{\mathrm{ens}}=60$ times to produce ensemble statistics, except for the case 1.C.2 which was performed $20$ times. In case 1.B, only vertically orientated bars of the active grid were actuated. For case 1.A, the grid was stationary with the wings aligned with the flow in the position of least blockage (see also Appendix A). The experimental conditions are listed in table 1.
@@
for $0\leq t\leq T_{\mathrm{wm}}$, where $S_{0}$ is the peak stroke, $\tau_{\mathrm{wm}}=6$ s is the group width in time as applied to the wavemaker signals and $T_{\mathrm{wm}}=24\,\mathrm{s}$ was the duration over which the wavemaker plungers were actuated. The surface elevation for one wave group measured at the SPIV measurement location is shown in figure 2(a).
Figure 2: (a) Example surface elevation of a single wave group from Experiment 1 as a function of time, measured by a wave probe at the measurement location. (b) Example from experiment 1 of an ensemble-average group surface elevation amplitude envelope as a function of time normalised by the measured group temporal width $\tau$ for case 1.D(see (4)). The time intervals for SPIV measurement (1-3) are shown with vertical dashed lines. (c) Surface elevation measurements of one ensemble from experiment 3 (cases 3.A and 3.B), which shows the onset of a regular wave train. The red box indicates the interval used for analysis.
@@ 2.1.2 Experiment 2: Regular waves
Cases 2.A and 2.B in tables 13 are from Experiment 2. Regular waves were generated propagating upstream atop two different flows with comparable mean velocity but different levels of turbulence as controlled by the active grid. A planar PIV set-up with a light sheet, orientated in the streamwise–vertical ($xz$) plane, measured the in-plane streamwise and vertical velocity components. The field of view was centred $L_{\mathrm{FOV}}=8.5$ m downstream of the surface plate’s trailing edge (see figure 1). Three cameras stacked vertically (from the top to bottom: two 16 megapixel cameras and one 5.5 megapixel camera) covered a field of view extending over the entire water depth ($h=0.80$ m) of the channel, roughly $21.7$ cm wide. The parameters for Experiment 2 are listed in table 1. For both flow cases, waves of two frequencies $0.94$ and $1.16$ Hz each with three different steepness values were generated. For each case, $2000$ PIV images were acquired at a sampling frequency of $0.86$ Hz; thus the total measurement time $T_{\mathrm{PIV}}$ was approximately $39$ min. Using a final pass with $48\times 48$ pixel interrogation window and a 50% overlap resulted in a velocity vector spacing of approximately $1.6$ mm. For both cases 2.A and 2.B, PIV measurements were also performed without wave generation to characterise ambient flow conditions, reported in Table 2.
@@ 2.1.3 Experiment 3: Onset of regular waves
Cases 3.A and 3.B in tables 13 are from Experiment 3. This set of measurements was taken at the highest acquisition frequency in our study ($15$ Hz) during a $25$ s interval just after the first arrival of regular waves as indicated in figure 2(c). The water depth was $0.50$ m, and a slower mean flow of $U_{0}=0.19$ m/s was used compared with the other cases (see table 1). The two cases are separated by the different sequences for the active grid. Case 3.A is similar to case 1.B as only vertically orientated grid bars were rotated at a random, top-hat-distributed rotation speed. Similarly for case 3.B, only the vertical grid bars were actuated, made to flap over an angle of $\pm 60^{\circ}$ about the fully open position during random time intervals of a top-hat-distributed duration between $0.5$ and $1.0$ seconds. (A full list of experimental conditions is found in appendix A.) No surface plate downstream of the active grid was used in this experiment, and the field of view in the streamwise–vertical plane (see figure 1c) was centred a distance $L_{\mathrm{FOV}}=8.5$ m downstream of the grid.
The wavemaker was actuated at $f_{0}=1.4$ Hz. Here, two steepnesses for each flow case are investigated, yielding four different steepness–turbulence combinations. For each of the four combinations, $32$ ensembles are measured, where the wave probe measurements of one ensemble are shown in figure 2(c). In order to avoid wave breaking at the leading edge of the wave train, the wavemaker stroke is linearly ramped up to the set amplitude over a period of $6$ seconds.
For the PIV measurements, a single 25-megapixel camera was used to cover the entire water column. Using a $64\times 64$ pixel interrogation window and a 50% overlap resulted in a final velocity vector spacing of approximately $3$ mm. The field of view is $0.45\,\mathrm{m}\times 0.5$ m. PIV images were acquired at $f_{\mathrm{ac}}=15$ Hz. Also here, PIV measurements were performed without wave generation to characterise the ambient flow conditions, reported in table 2. For these background measurements $2000$ snapshots were captured at $1.0$ Hz.
@@ 3.1 Flow and wave characteristics
The measured physical characteristics of mean flow, turbulence and waves are listed in table 2, and some derived quantities we will make use of are quoted in table 3 for convenience.
Case $U_{0}$ $k_{0}$ $\tau$ $u_{\mathrm{rms}},v_{\mathrm{rms}},w_{\mathrm{rms}}]$ $e$ $L_{x}^{x}$ Steepness
$(\mathrm{m/s})$ $(\mathrm{rad/m})$ $(\mathrm{s})$ $(\mathrm{cm/s})$ $(\mathrm{cm^{2}/s^{2}})$ $(\mathrm{cm})$
1.A 0.34 9.5 7.3 $[0.71,0.68,0.58]$ 0.65 $5.1$ $0.20$
1.B 0.33 9.2 7.2 $[1.1,1.0,0.73]$ 1.3 $26$ $0.20$
1.C 0.33 8.9,9.0 7.9 $[1.6,1.2,0.92]$ 2.4 $32$ $0.15,0.22$
1.D 0.34 9.3 7.8 $[1.7,1.7,1.3]$ 3.7 2020 $0.22$
2.A.1 0.30 6.1 8.3 [2.5,,1.8][2.5,-,1.8] 6.4 - $0.09,0.14,0.18$
2.A.2 0.30 12.1 - [2.5,,1.8][2.5,-,1.8] 6.4 - $0.15,0.19,0.21$
2.B.1 0.30 6.1 - [1.6,,1.2][1.6,-,1.2] 2.7 - $0.07,0.12,0.17$
2.B.2 0.30 12.0 - [1.6,,1.2][1.6,-,1.2] 2.7 - $0.14,0.17,0.21$
3.A 0.19 12.9 - [1.2,,0.84][1.2,-,0.84] 1.4 5.4 $0.11,0.18$
3.B 0.19 13.0 - [0.87,,0.70][0.87,-,0.70] 0.87 5.1 $0.11,0.18$
Table 2: Measured flow quantities: mean current $U_{0}$, carrier-wave number $k_{0}$, measured group temporal width in the lab frame $\tau$, root-mean-square turbulent velocity components (subscript ‘rms’), turbulent kinetic energy $e$ (calculated for Experiments 2 and 3 as discussed in main text), and streamwise–streamwise integral scale $L_{x}^{x}$. Steepness is given as $k_{0}a_{\mathrm{p}}$ for Experiment one where $a_{\mathrm{p}}$ is peak amplitude, and $ak_{0}$ in Experiments 2 and 3. Where several values of steepness are listed (cases 1.C and 2.A3.B), these are referred to elsewhere as 1.C.1, 1.C.2, 2.A.1.1, 2.A.1.2, etc.
Case $\lambda_{0}$ $c(k_{0})$ $u_{s}(0)$ $T_{0}$ $\tau_{0}$ $|u_{\mathrm{rf}}|$
@@
1.A $0.68$ $1.0$ $4.1$ $0.65$ 2.4 -
1.B $0.70$ $1.0$ $4.2$ $0.66$ 2.6 -
1.C $0.71,0.70$ $1.0$ $2.3,5.1$ $0.67$ 2.9 -
1.D $0.68$ $1.0$ $5.0$ $0.66$ 2.8 -
2.A.1 $1.03$ $1.3$ $1.0,2.5,4.1$ $0.81$ - $1.1,2.5,4.2$
2.A.2 $0.52$ $0.90$ $2.0,3.3,4.0$ $0.58$ - $1.0,1.7,2.1$
2.B.1 $1.03$ $1.3$ $0.62,1.8,3.7$ $0.81$ - $0.6,1.9,3.8$
2.B.2 $0.52$ $0.90$ $1.8,2.6,4.0$ $0.58$ - $0.9,1.4,2.1$
3.A $0.49$ $0.87$ 1.1, 2.8 0.56 - $0.8,2.2$
3.B $0.48$ $0.87$ 1.1, 2.8 0.56 - $0.8,2.2$
Table 3: Wave quantities derived from measured values in Table 2. Here $\lambda_{0}=2\pi/k_{0}$ is the carrier wavelength, $c(k_{0})=\sqrt{g/k_{0}}$ is the carrier-wave phase velocity, $u_{s}(0)=(ak_{0})^{2}c(k_{0})$ is the Stokes drift velocity of the carrier wave at the surface, $T_{0}=\lambda_{0}/c(k_{0})$ is the intrinsic wave period, $\tau_{0}$ the intrinsic group width from Eq.0(5), and $u_{\mathrm{rf}}$ is the Eulerian return flow from Eq.0(8).
An assumption of deep water is well satisfied since in all our cases $k_{0}h\gtrsim 3.6$ hence $1>\sqrt{\tanh(k_{0}h)}>0.999$. Here and henceforth, $U_{0}$ is the mean absolute surface speed in the absence of waves (i.e., the mean surface velocity is $\mathbf{U}_{0}=(-U_{0},0,0)$) and in all our cases the velocity profile is sufficiently depth-uniform that vertical shear affects wave dispersion negligibly (e.g., Ellingsen & Li, 2017). The wavenumber $k_{0}$ and the wavemaker carrier frequency $f_{0}$ are related approximately by $2\pi f_{0}=\sqrt{gk_{0}}-U_{0}k_{0}$, or equivalently $f_{0}=(c(k_{0})-U_{0})/\lambda_{0}$ with $c(k_{0})$ from Eq. (2) and the wavelength is $\lambda_{0}=2\pi/k_{0}$. We will use the measured (as opposed to derived) value of $\lambda_{0}$ and the wavenumber $k_{0}$ accordingly. Values for these quantities are listed in tables 2 and 3, respectively. The time it takes for the waves to propagate from wave-maker to the edge of the surface plate is around $40$ s for Experiment 1, $25$ s for Cases $\textit{2.A}.1$ and $\textit{2.B}.1$, $56$ s for Cases $\textit{2.A}.2$ and $\textit{2.B}.2$ and $35$ s for Experiment 3.
The root-mean-square (r.m.s.) of the turbulent velocity fluctuation after subtracting the average is defined as ${\boldsymbol{u}}_{\mathrm{rms}}=[u_{\mathrm{rms}},v_{\mathrm{rms}},w_{\mathrm{rms}}]$ representing the streamwise, spanwise, and vertical components, respectively. The turbulent kinetic energy is as $e=\frac{1}{2}|{\boldsymbol{u}}_{\mathrm{rms}}|^{2}$. In cases 2.A3.B, only $u_{\mathrm{rms}}$ and $w_{\mathrm{rms}}$ are available, so we use instead $e=\frac{1}{2}(u_{\mathrm{rms}}^{2}+2w_{\mathrm{rms}}^{2})$, an assumption of cross-plane isotropy, as is the standard convention for grid turbulence (e.g., Comte-Bellot & Corrsin, 1966; Lavoie et al., 2005) and was observed for our facility by Jooss et al. (2021). It is not quite satisfied for our flow as Table 2 shows for Experiment 1. However, given the inevitable anisotropy of the flow as the free surface is approached (also with no waves) and consequent depth variation of $e$ closer to the surface than approximately one integral scale (the blockage layer, see, e.g. Aarnes et al., 2025), representing turbulent kinetic energy (TKE) by a single number is always a simplified picture. The precise value of $e$ does not significantly affect our conclusions, and given that no simple expression for $e$ in terms of $u_{\mathrm{rms}}$ and $w_{\mathrm{rms}}$ can reconcile all cases in Experiment 1, we choose to go with the conventional expression.
The mean velocity $U_{0}$ was calculated from averaging all streamwise velocities over the lower part of the field of view; averaging was performed over 12-12 cm <z<5<z<-5 cm for Experiment 1 and 20-20 cm <z<10<z<-10 cm for Experiments 2 and 3. Full profiles of the Eulerian-mean velocity profiles for all experimental cases may be found in the Supplementary Materials.